OpenStudy (anonymous):

A truck covers 40.0 m in 7.10 s while smoothly slowing down to a final velocity of 2.75 m/s. (a) Find the truck's original speed.

OpenStudy (anonymous):

To solve this use the kinematic equations. You are given distance (d), time (t), and final velocity (Vf). To solve this kind of problem take a look at your 3 kinematic equations and get any variable you can with what you are given. The 3 equations are Vf=Vo + at (final velocity= inital velocity + acceleration x time), ∆d=Voxt+1/2at^2, and Vf^2= Vo^2 +2at. for this specific problem you will use two equations to set up a systems of equations. First we know 40=Vox7.1 + 1/2 a7.1^2, and we also know 2.75= Vo + a7.1. So a=2.75-Vo/7.1. Then plug this back into the first equation, and youll get 40=7.1Vo + 25.205(2.75-Vo/7.5). you cannow solve for Vo which will be equal too8.22 m/s. Hope this helps <(^_^)>

OpenStudy (anonymous):

v = 2.75m/s , s = 40 m , t=7.10 s , by using two equation of motion 1) v=u+at 2.75 = u + 7.10a 2) s= ut +1/2 at^2 40 = 7.10u + 1/2 (7.10^2)a we can get the value of u and a u will be the original speed